40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 336     Entry time: Fri Feb 22 15:16:33 2008
Author: Andrey 
Type: Update 
Category: PEM 
Subject: ITMX Accelerometer is NOT broken 

As I wrote in message 330, there was a bad signal from ITMX accelerometer. I have found the reason: the BNC-cable which goes from the black board with switches for accelerometer gain (1,10,100) towards DAQ-tower was completely disconnected from that black board with gain-switches. The end of the long BNC-cable was on the floor. Therefore, it was totally impossible to see any accelerometer signal. The cable that I am writing about should transport the signal from ITMX_X accelerometer.

Now all the BNC-connections seem to be in good shape, and spectra of accelerometers near ITMX and ETMX , both of them are in x-directions, are very much similar.
Attachment 1: Accelerom-ITMX-Feb23.jpg  71 kB  | Show | Hide all | Show all
Attachment 2: Accelerom-ETMX-Feb23.jpg  71 kB  | Hide | Hide all | Show all
Accelerom-ETMX-Feb23.jpg
ELOG V3.1.3-